27. Find the sum of the sequence 14 (k2 – 4) k=1
Answer #1

내 EK) 셔 (4)+(24) (34) + (424) + (==4) + =) +H4]+€=4)+A=4) +64| 44424) 43르니 거 19 2. K니 959 셔